# Claim Check-Worthiness Detection: How Well do LLMs Grasp Annotation Guidelines?

### v1

@@ Abstract
The increasing threat of disinformation calls for automating parts of the fact-checking pipeline. Identifying text segments requiring fact-checking is known as claim detection (CD) and claim check-worthiness detection (CW), the latter incorporating complex domain-specific criteria of worthiness and often framed as a ranking task. Zero- and few-shot LLM prompting is an attractive option for both tasks, as it bypasses the need for labeled datasets and allows verbalized claim and worthiness criteria to be directly used for prompting. We evaluate the LLMs’ predictive and calibration accuracy on five CD/CW datasets from diverse domains, each utilizing a different worthiness criterion. We investigate two key aspects: (1) how best to distill factuality and worthiness criteria into a prompt and (2) what amount of context to provide for each claim. To this end, we experiment with varying the level of prompt verbosity and the amount of contextual information provided to the model. Our results show that optimal prompt verbosity is domain-dependent, adding context does not improve performance, and confidence scores can be directly used to produce reliable check-worthiness rankings.
@@
Laura Majer  and Jan Šnajder TakeLab, Faculty of Electrical Engineering and Computing, University of Zagreb {laura.majer, jan.snajder}@fer.hr
@@ 1 Introduction
Automating fact-checking is becoming crucial in response to rising amounts of data and disinformation. Fact-checking is typically done on claims – to warrant fact-checking, a claim must be both factual (i.e., related to purported facts) and check-worthy (i.e., of interest to society). The NLP tasks of identifying factual and check-worthy claims are known as claim detection (CD) and claim check-worthiness detection (CW), respectively. While both tasks are typically defined as classification tasks, CW can also be framed as a ranking task, mimicking the prioritization process employed by fact-checking organizations (FullFact, 2020).
@@
Both CD and CW are challenging for several reasons. Firstly, the underlying concepts of factual claims and worthiness resist straightforward definitions. To grasp factuality, Konstantinovskiy et al. (2021) presented a thorough categorization of factual claims, while Ni et al. (2024) provided a definition distinguishing opinions. Defining check-worthiness is made more challenging by its subjective, context-dependent nature and temporal variability. Assessing it usually requires choosing more specific criteria, such as relevance to the general public (Hassan et al., 2017a) or policymakers, potential harm (Nakov et al., 2022), or alignment with a particular topic (Stammbach et al., 2023; Gangi Reddy et al., 2022)). Another challenge is identifying the situational context (including previous discourse and speaker information) required to determine claim factuality and check-worthiness.
@@
CD and CW have been addressed using traditional supervised ML and fine-tuning pre-trained language models. However, acquiring labeled datasets can be challenging – they must align with a specific language, domain, and genre and match the desired factuality and worthiness criteria. Moreover, dataset annotation is costly and requires redoing if criteria change. LLMs present a viable alternative to supervised methods owing to their strong zero- and few-shot performance (Kojima et al., 2022; Brown et al., 2020). Over time, fact-checking organizations have refined principles for claim prioritization, and few-shot prompting offers a seamless way to transfer this knowledge to the model. Thus, an effective strategy might entail zero- and few-shot prompting with check-worthiness criteria from annotation guidelines.
@@
In this paper, we study the predictive and calibration accuracy of zero- and few-shot LLM prompting for CD and CW. We experiment with five datasets, each with a different factuality or worthiness criterion outlined in the accompanying annotation guidelines. We investigate two key aspects: (1) how to best distill factuality and worthiness criteria from the annotation guidelines into the prompt and (2) what amount of context to provide for each claim. For (1), we experiment with varying the level of prompt verbosity, starting from brief zero-shot prompts to more detailed few-shot prompts that include examples. For (2), we experiment with expanding the prompt with co-text and other components of the claim’s situational context. Finally, we consider CW as a ranking task using LLM’s confidence scores as a proxy for priority. We show that gpt-4-turbo with worthiness criteria adopted from annotation guidelines can yield accuracy and ranking scores comparable to or outperforming existing CD/CW methods. Our results demonstrate the potential of using LLMs for claim check-worthiness detection with minimal prompt engineering.
@@ 2 Related Work
The CD and CW tasks constitute the first part of the fact-checking pipeline. Typically framed as classification tasks, they are handled using traditional supervised machine learning (Hassan et al., 2017b; Wright and Augenstein, 2020; Hassan et al., 2017a; Gencheva et al., 2017), or fine-tuning PLMs (Stammbach et al., 2023; Sheikhi et al., 2023).
Recently, the use of LLMs for CD and CW is starting to take on. Sawinski et al. (2023) and Hyben et al. (2023) compare the performance of fine-tuned PLMs with LLMs using zero- and few-shot learning as well as fine-tuning. Although zero- and few-shot approaches for LLMs underperform, the authors note their reliance on internal definitions of worthiness and limited prompt testing. As part of the fully automated fact-checking system relying only on LLMs, Li et al. (2023) implement a CD module using a verbose few-shot prompt, yet they do not report performance metrics. Finally, Ni et al. (2024) tackle CD by proposing a three-step prompting approach to examine model consistency. However, neither Li et al. (2023) nor Ni et al. (2024) address the CW task. To our knowledge, there is no work on CW focused on describing specific worthiness criteria using verbose prompts.
@@ 3 Datasets
Our experiments use five datasets in English covering diverse topics and genres:
@@
# instances used 1,032 251 570 6,129 816
@@
Co-text 4 previous, on request - not available inconclusive 1 previous, 1 following
Agreement -* 0.75/0.7 0.47 0.405 0.69
Agreement metric - Fleiss-$\kappa$ Krippendorff-$\alpha$ Krippendorff-$\kappa$ Cohen-$\kappa$
Table 1: Characteristics of the CD and CW datasets used in our experiments. *CB reported no agreement evaluation, but the test set used is agreed upon by experts.
ClaimBuster (CB)
(Hassan et al., 2017a) is a widely used dataset of claims from USA presidential debates. It uses ternary labels (non-factual, unimportant factual, check-worthy factual) and thus allows for the distinction between check-worthy and unimportant factual claims, therefore covering both the CD and CW tasks;
CLEF CheckThat!Lab 2022 (CLEF)
(Alam et al., 2021) is a dataset of tweets relating to COVID-19. It comprises two sets: a set of tweets containing claims and a subset of those containing check-worthy claims, thus covering both the CD and CW tasks. Check-worthiness is defined as the need for professional fact-checking, excluding claims that are “too trivial to check”;
EnvironmentalClaims (ENV)
(Stammbach et al., 2023) focuses on environmental articles and reports. It defines specific criteria for an environmental claim that extend beyond the topic itself (e.g., the claim should be explicit, focus on environmental impact);
NewsClaims (NEWS)
(Gangi Reddy et al., 2022) is a dataset of sentences from news articles on COVID-19, with metadata available for positives (speaker, object, claim span). The annotators were asked to judge whether a claim falls into one of the four topic-specific categories – essentially forming the worthiness criteria;
PoliClaim (POLI)
(Ni et al., 2024) covers the same topic as ClaimBuster (politics, speeches of governors) but labels only factual claims, leaving out check-worthiness. The binary labels are obtained by aggregating responses to two questions.
We selected these datasets because they disclose annotation guidelines to some degree of detail and cover different topics, genres, and worthiness criteria. Table 1 summarizes the datasets’ characteristics (cf. Appendix A for more details). The CB and CLEF datasets cover both tasks, where CB uses ternary labels annotated at once, while CLEF uses binary labels and separate annotation questions for CD and CW. The datasets were either originally annotated using a binary scheme (ENV), Likert scale (CLEF-CW), multi-class based on topic (NEWS), or a follow-up prompt for uncertain instances (POLI). All the datasets provided aggregated binary labels, except CB, where aggregation from ternary into binary CD and CW labels is straightforward. The reported inter-annotator agreement for POLI and CLEF is substantial (Landis and Koch, 1977), while the agreement for ENV and NEWS datasets is moderate, confirming the complexity of the domain-dependant CW task.
4 Experiments and Results
In our experiments, we use OpenAI models gpt-turbo-3.5 and gpt-4-turbo (cf. Appendix C for a description of parameters). We also experimented with smaller, open-source models, but their responses often did not match the target labels.
@@ 4.1 Prompt verbosity
We first investigate how prompt verbosity affects LLMs’ predictive accuracy. We hypothesize that the optimal verbosity level depends on the dataset, reflecting the factuality and worthiness criteria differences between the domains. While a brief prompt might lack essential details, a comprehensive prompt featuring extensive definitions and examples may make the task more ambiguous for the model. Based on the content and style of annotation guidelines, we define the following four levels of verbosity (cf. Appendix E for full prompts):
@@
serves as the baseline. We use a naive zero-shot prompt relying on the models’ internal definitions of factual and check-worthy for the CD and CW tasks, respectively. As this prompt does not include the specific worthiness criteria from the guidelines, it serves as a domain-agnostic baseline;
@@
uses prompts that include the task definition and the set of possible labels but omit detailed explanations of the labels or principles;
@@
expands on V1 by adding a more detailed explanation of the labels or general annotation principles (or both, in the case of PoliClaim);
@@
builds on V2 with examples provided to annotators. This level is closest to annotation guidelines, encompassing all or most information that the datasets’ authors provided in the papers accompanying the datasets.
Table 2 shows the F1 scores by verbosity level for gpt-4-turbo (gpt-3.5-turbo generally performed worse; cf. Appendix A.2). Predictive accuracy for CD is generally higher than for CW, proving that capturing worthiness is more difficult than identifying factual claims. The optimal verbosity level is not consistent across datasets: the performance increases with verbosity levels for CB, but the trend is reversed for ENV. We observe no consistent trend for CLEF, POLI, and NEWS datasets. The naive baseline prompt outperforms our prompts on POLI and NEWS datasets. On the other hand, our prompts outperform some of the previously reported results (cf. Appendix A.2), proving the potential of prompting with annotation guidelines.

### v2

@@ Abstract
The rising threat of disinformation underscores the need to fully or partially automate the fact-checking process. Identifying text segments requiring fact-checking is known as claim detection (CD) and claim check-worthiness detection (CW), the latter incorporating complex domain-specific criteria of worthiness and often framed as a ranking task. Zero- and few-shot LLM prompting is an attractive option for both tasks, as it bypasses the need for labeled datasets and allows verbalized claim and worthiness criteria to be directly used for prompting. We evaluate the LLMs’ predictive accuracy on five CD/CW datasets from diverse domains, using corresponding annotation guidelines in prompts. We examine two key aspects: (1) how to best distill factuality and worthiness criteria into a prompt, and (2) how much context to provide for each claim. To this end, we experiment with different levels of prompt verbosity and varying amounts of contextual information given to the model. We additionally evaluate the top-performing models with ranking metrics, resembling prioritization done by fact-checkers. Our results show that optimal prompt verbosity varies, meta-data alone adds more performance boost than co-text, and confidence scores can be directly used to produce reliable check-worthiness rankings.
@@
Laura Majer  and Jan Šnajder Text Analysis and Knowledge Engineering Lab University of Zagreb, Faculty of Electrical Engineering and Computing {laura.majer, jan.snajder}@fer.hr
@@ 1 Introduction
Figure 1: Using annotation guidelines, we craft zero- and few-shot LLM prompts for claim and claim check-worthiness detection, varying the level of prompt verbosity and the amount of provided context. We evaluate the LLMs using classification and ranking metrics.
The global spread of information, coupled with mis- and disinformation, is increasing the demand for fact-checking (News, 2022; Idrizi and Hanafin, 2023), highlighting the need for automation. However, complete automation may not be ideal; for instance, PolitiFact, which used ChatGPT to verify previously fact-checked claims, faced issues like inconsistency, knowledge limitations, and misleading confidence (Abels, 2023). Nevertheless, they see potential in language models for assisting fact-checkers, particularly in identifying claims worth verifying. Similarly, FullFact has highlighted the lack of effective claim selection tools as a major workflow challenge (FullFact, 2020).
@@
To warrant fact-checking, a claim must be both factual (i.e., related to purported facts) and check-worthy (i.e., of interest to society). The NLP tasks of identifying factual and check-worthy claims are known as claim detection (CD) and claim check-worthiness detection (CW), respectively. The tasks make up the first component of the automatic fact-checking pipeline. While both are typically defined as classification tasks, CW can also be framed as a ranking task, mimicking the prioritization process employed by fact-checking organizations.
@@
Both CD and CW are challenging for several reasons. Firstly, the underlying concepts of factual claims and check-worthiness resist straightforward definitions. To grasp factuality, Konstantinovskiy et al. (2021) presented a thorough categorization of factual claims, while Ni et al. (2024) provided a definition distinguishing opinions. Regardless of these variations, factual could be deemed universal and self-explanatory, unlike check-worthiness, a term frequently used in previous research. Defining check-worthiness is made more challenging by its subjective, context-dependent nature and temporal variability. Assessing it usually requires choosing more specific criteria, such as relevance to the general public (Hassan et al., 2017a) or policymakers, potential harm (Nakov et al., 2022), or alignment with a particular topic (Stammbach et al., 2023; Gangi Reddy et al., 2022)). Another challenge is identifying the situational context (including previous discourse and speaker information) required to determine claim factuality and check-worthiness. For example, in CW annotation campaigns (Hassan et al., 2017a; Gangi Reddy et al., 2022), annotators are typically presented with surrounding sentences to aid their assessment.
@@
The CD and CW tasks have been approached using both traditional supervised machine learning and fine-tuning pre-trained language models, both of which depend on labeled data. However, obtaining such datasets can be challenging as they need to align with specific languages, domains, and genres and meet desired factuality and worthiness criteria. Moreover, dataset annotation is costly and requires redoing if criteria change. LLMs present a viable alternative to supervised methods owing to their strong zero- and few-shot performance (Kojima et al., 2022; Brown et al., 2020). Over time, fact-checking organizations have refined principles for claim prioritization, and zero- and few-shot prompting offers a seamless way to transfer this knowledge to the model. Thus, an effective strategy might entail zero- and few-shot prompting with check-worthiness criteria from annotation guidelines. The challenge, however, is that LLMs often exhibit sensitivity to variations in prompts (Mizrahi et al., 2024) and unreliability (Si et al., 2023).
@@
In this paper, we study the predictive and calibration accuracy of zero- and few-shot LLM prompting for CD and CW. We experiment with five datasets, each with a different factuality or worthiness criterion outlined in the accompanying annotation guidelines. We investigate two key aspects: (1) how to best distill factuality and worthiness criteria from the annotation guidelines into the prompt and (2) what amount of context to provide for each claim. For (1), we experiment with varying the level of prompt verbosity, starting from brief zero-shot prompts to more detailed few-shot prompts that include examples. For (2), we expand the prompt with co-text and other components of the claim’s situational context. Furthermore, inspired by the fact-checker’s prioritization process, we consider CW as a ranking task, using LLM confidence scores as a proxy for determining priority. Figure  1 depicts the workflow of our experiments. We show that prompting with worthiness criteria adopted from annotation guidelines can yield accuracy and ranking scores comparable to or surpassing existing CD/CW methods. Although optimal prompt verbosity varies across datasets, certain in-domain trends can be observed across models. We also find that the impact of adding context is greater for lower verbosity levels, while meta-data is more beneficial than co-text. Finally, we show that confidence scores can be directly used to produce reliable check-worthiness rankings.
Our contributions include analyzing LLM performance in terms of (1) prompt detail, (2) provided context, and (3) variations across domains and worthiness criteria.
@@ 2 Related Work
Developing a fully automated fact-checking system is appealing for both its applicability and the challenge it presents (Hassan et al., 2017c; Li et al., 2023). However, Glockner et al. (2022) question the purpose of such a system, pointing to its reliance on counter-evidence that may not be available for newly coined disinformation. This motivates a shift toward human-in-the-loop approaches and automating parts of the fact-checking pipeline.
The CD and CW tasks constitute the first part of the fact-checking pipeline and are meant to select parts of the input for which fact-checking is possible (CD) or deemed necessary (CW). Typically framed as classification tasks, the CD and CW tasks are handled using traditional supervised machine learning (Hassan et al., 2017b; Wright and Augenstein, 2020; Hassan et al., 2017a; Gencheva et al., 2017) or fine-tuning pre-trained language models (Stammbach et al., 2023; Sheikhi et al., 2023). Methods of solving include rich sentence and context-level features (Gencheva et al., 2017), speaker, object, and claim span identification (Gangi Reddy et al., 2022), or incorporating domain-specific knowledge by combining ontology and sentence embeddings (Hüsünbeyi and Scheffler, 2024). CW can also be framed as a ranking task (Jaradat et al., 2018; Gencheva et al., 2017), mimicking the prioritization of claims by fact-checking organizations.
Recently, the use of LLMs for CD and CW is starting to take on. Sawinski et al. (2023) and Hyben et al. (2023) compare the performance of fine-tuned language models such as BERT with LLMs using zero- and few-shot learning as well as fine-tuning. Although zero- and few-shot approaches for LLMs underperform, the authors note their reliance on internal definitions of worthiness and limited prompt testing. As part of the fully automated fact-checking system relying only on LLMs, Li et al. (2023) implement a CD module using a verbose few-shot prompt, yet they do not report performance metrics. Finally, Ni et al. (2024) tackle CD by proposing a three-step prompting approach to examine model consistency. However, neither Li et al. (2023) nor Ni et al. (2024) address the CW task. To our knowledge, there is no work on CW focused on describing specific worthiness criteria using verbose prompts.
@@ 3 Datasets
Our experiments utilize five datasets in English covering diverse topics and genres. Examples from each dataset are presented in Table 1. We next describe each dataset in more detail, including the CD and CW criteria used.
Dataset Label Example CB I would do the opposite in every respect. O I have met with the heads of government bilaterally as well as multilaterally. Fifty percent of small business income taxes are paid by small businesses. CLEF If the vaccine was dangerous they would’ve given it to poor people first, not politicians and billionaires. O Today, FDA approved the first COVID-19 vaccine for the prevention of #COVID19 disease in individuals 16 years of age and older. They said the vaccine stopped transmission. Now they are lying and saying they didn’t. Video proof here ENV We Love Green! The environment is at the heart of Parisian electro-pop music festival We Love Green. All pension fund clients have a target for carbon reduction of the equity investments. NEWS In Germany, RT has also amplified voices questioning the threat of COVID-19, and calling testing and mask-wearing into question. "If you wash and dry a cloth face mask on high heat, then you should be good to go," according to professor Travis Glenn. POLI As I have said all along, the courts are where we will win this battle. O I promised that our roads would be the envy of the nation.
Table 1: Examples from the datasets used. ✗= non-factual claim, O = factual claim, ✓= check-worthy claim
ClaimBuster (CB)
(Hassan et al., 2017a) is a widely used dataset of claims from USA presidential debates, featuring ternary labels (non-factual, unimportant factual, check-worthy factual) that distinguish between check-worthy and unimportant factual claims. This setup addresses both the CD and CW tasks. Claims are deemed check-worthy if the general public would be interested in their veracity. However, no specific definition of factuality is provided – unimportant factual claims are defined as those lacking check-worthiness.
CLEF CheckThat!Lab 2022 (CLEF)
(Alam et al., 2021) contains tweets about COVID-19, with two parts: a set of tweets with claims and a subset with check-worthy claims, addressing both CD and CW tasks. Check-worthiness is defined as the need for professional fact-checking, excluding jokes, trivial claims, or those deemed uninteresting. Factual claims are defined as sentences that assert something is true and can be verified using factual information, such as statistical data, specific examples, or personal testimony.
EnvironmentalClaims (ENV)
(Stammbach et al., 2023) is compiled from environmental articles and reports. The dataset focuses on check-worthy environmental claims related to green-washing in marketing strategies. The authors defined specific criteria for an environmental claim that extend beyond the topic itself (e.g., highlighting the positive environmental impact of a product, not being too technical). The annotators were instructed to label only the explicit claims, discouraging the selection of claims with inter-sentence coreferences.
NewsClaims (NEWS)
(Gangi Reddy et al., 2022) comprises sentences from news articles on COVID-19, with metadata available for positives (speaker, object, claim span). The annotators were asked to judge whether a claim falls into one of the four topic-specific categories, which essentially formed the worthiness criteria, even though check-worthiness was not explicitly mentioned in the guidelines. The dataset includes both check-worthy and non-check-worthy claims with inter-sentence coreferences (e.g., That’s also false), which typically require inspecting the surrounding context to determine their check-worthiness (we estimate this applies to about 10% of claims in the test set).
PoliClaim (POLI)
(Ni et al., 2024) covers the same topic as ClaimBuster (politics, speeches of governors) but labels only verifiable claims, leaving out check-worthiness. The authors provided detailed guidelines on verifiable claims, emphasizing the need for specificity and differentiation from opinions lacking factual basis. To handle ambiguous cases, they employed a ternary (Yes, No, Maybe) annotation scheme. Maybe indicates that a claim may contain factual information but does not fully meet all criteria. For claims labeled as Maybe, annotators answered a follow-up Yes-No question to determine whether the claim leans toward factual information or subjective opinion. As with NEWS, inter-sentence coreference was considered; since the claims are extracted from political speeches, many of them include personal pronouns (I, we), which necessitates coreference resolution to identify the claimant or subject.
@@
# instances used 1,032 251 275 1,622 816
label distribution 731/238/63 102/110/39 198/67 811/811 295/521
@@
Co-text 4 preceding, on request not available inconclusive 1 preceding, 1 following
Agreement –* 0.75/0.7 0.47 0.405 0.69
Agreement metric Fleiss-$\kappa$ Krippendorff-$\alpha$ Krippendorff-$\kappa$ Cohen-$\kappa$
Table 2: Characteristics of the CD and CW datasets used in our experiments. *CB reported no agreement evaluation, but the test set used is agreed upon by experts. Label distribution in order: ✗/ O / ✓
We use these five datasets because they provide detailed annotation guidelines and cover various topics, genres, and worthiness criteria. Table 2 summarizes their characteristics (see Appendix A for details). The CB and CLEF datasets address both CD and CW tasks, with CB using ternary labels annotated together and CLEF using binary labels with separate questions for CD and CW. The five datasets were originally annotated using a binary scheme (ENV), Likert scale (CLEF-CW), multi-class (NEWS), or a follow-up prompt for uncertain instances (POLI). All datasets have aggregated binary labels, except CB, where aggregation from ternary into binary labels is straightforward. The reported inter-annotator agreement is substantial for POLI and CLEF (Landis and Koch, 1977), but moderate for ENV and NEWS, reflecting the complexity of the domain-dependent CW task.
4 Experimental Setup
In our experiments, we use both closed-source and open-source LLMs. For closed source, we use OpenAI models gpt-turbo-3.5 and gpt-4-turbo. For open-source models, due to hardware constraints, we chose Llama 3 8B Instruct, which is the top performer in its parameter class. To ensure reproducibility and encourage deterministic behaviour, we prompt GPT models with the temperature setting of 0 along with a fixed seed parameter and use greedy sampling with top_p=1 for open-source models. We also experimented with other open-source models. Mistral 7B Instruct v0.2 was not compliant with the provided labels, instead giving open-ended answers, even for less verbose prompts. See Appendix B for more detailed information on models.
@@ 4.1 Prompt Verbosity
We first investigate how prompt verbosity affects LLMs’ predictive accuracy. We hypothesize that the optimal verbosity level depends on the dataset, reflecting the factuality and worthiness criteria differences between the domains. While a brief prompt might lack essential details, a comprehensive prompt featuring extensive definitions and examples may make the task more difficult to solve. Across datasets and for each prompt level, we aim to preserve the original wording and typography of the annotation guidelines as much as possible since we aim to establish whether guidelines without much intervention can be used as prompts for up-to-par performance. We additionally instruct the model to reply using only the provided labels without additional explanation to increase compliance and streamline evaluation. For POLI, we use the same question structure as in the annotation – for instances where the model responded with Maybe, we prompt it again with the follow-up question, providing previous responses in the prompt.
Based on the content and style of annotation guidelines, we define the following four levels of verbosity (cf. Appendix D for full prompts for four verbosity levels across the five datasets):
@@
serves as the baseline. We use a naive zero-shot prompt, relying on internal definitions of the model. For the CD task (for the CB, CLEF and POLI datasets), we use the following prompt: “Does the following sentence/statement/tweet contain a factual claim? Answer only with Yes or No.” For the CW task (for the CB, CLEF, NEWS and ENV datasets) we use the following prompt: “Does the following sentence/statement/tweet contain a check-worthy claim? Answer only with Yes or No.” As these prompts do not include the specific factuality or worthiness criteria from the guidelines, they serve as a domain-agnostic baseline;
@@
uses prompts that include the task definition and the set of possible labels but omit detailed explanations of the labels or principles. For example, for the CB dataset, the three categories of non-factual, unimportant factual, and check-worthy factual sentences are introduced but not explained;
@@
expands on V1 by adding a more detailed explanation of the labels or general annotation principles (or both, in the case of PoliClaim). Some principles include avoiding implicit assumptions (ENV), defining check-worthiness criteria based on public interest (CB), and categorizing claims that non-professionals can verify as non-check-worthy (CLEF);
@@
builds on V2 by including examples from the original annotation guidelines. This level closely aligns with annotation guidelines, encompassing all or nearly all information the datasets’ authors provide in their accompanying papers. [^1] CB and ENV documented additional examples (typically 20–30 examples) provided to the annotators. We did not include these examples. The examples are provided either along with the labels (CB), separately in a few-shot fashion (ENV), or both (POLI).
4.2 Amount of Context
In real-world scenarios, claims are rarely evaluated in isolation. Accordingly, annotators working with CD and CW datasets were usually provided with some contextual information, consisting of the claim’s co-text and metadata. Regarding co-text, the quantity varied between datasets (cf. Table 2), as did its significance – sometimes it was provided as additional guidance (CB), while in other cases, it was deemed crucial for assigning labels (POLI, NEWS). For NEWS, the amount of provided co-text is inconclusive, so we decided to omit it from co-text expansion. This difference highlights that co-text is both another undefined aspect of CD and CW, and that it can vary across domains. Similarly, metadata such as speaker, affiliation, occasion, and date were revealed only during annotation for CLEF-CW and were not available in the dataset itself. However, metadata is available for the CB and POLI datasets, while NEWS provides metadata only for positives, making it unusable for our experiments. Adding metadata might lead to biases, yet it could offer essential information, depending on the worthiness criterion.
We investigate how LLMs’ predictive accuracy depends on the amount of situational context provided to the model. To this end, we leverage the context information available in the CB and POLI datasets and expand the prompts in three variants: